if a box has a volume of 900 cubic cm, and its width is 3cm and its height is 5 cm, what is its length

Answer by nyc_function(2626)

Here is the formula:
V = L*W*L
900 = L*3*5
900 = 15L
900/15 = L

60 = L
The length is 60cm.